4) What will be the output of the following Java code snipper? x=99, y=100; print x and y a. Error b· x=99 y= 100 x=100 y=99 c. d. x-99 y-99 5) osider the fellowing C code snippet:

please solve this fully
Answer #1

4) Answer: (c) x = 100 y = 99

x * y = 99 * 100 = 9900

x = y = 100 (y value will be stored in x)

y = (x * y) / (x = y)

=> 9900 / 100 = 99

Therefore, x = 100 y = 99